OpenStudy (sawq):

Urgent question! Will fan and medal! What is the value of z for the equation fraction 1/ 2z = −3/4 + 1/4z? −3 −1 1 −3 −1 1 −3 −1 1 −3 −1 1 3 a. -3 b. -1 c. 1 d. 3

OpenStudy (sawq):

woah sorry the options are: a. -3 b. -1 c. 1 d. 3

OpenStudy (sawq):

@dan815

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

is it \[\frac{1}{2}z=-\frac{3}{4}+\frac{1}{4}z\]\

OpenStudy (lynfran):

-3

OpenStudy (sawq):

ya @misty1212

OpenStudy (misty1212):

if you hate fractions like i do, multiply both sides by 4 and start with \[2z=-3+z\] then you can solve in two steps

OpenStudy (misty1212):

actually one step subtract \(z\) from both sides, that is all